# Failure Modes of LLMs for Causal Reasoning on Narratives

### v2

@@ Abstract
In this work, we investigate the causal reasoning abilities of large language models (LLMs) through the representative problem of inferring causal relationships from narratives. We find that even state-of-the-art language models rely on unreliable shortcuts, both in terms of the narrative presentation and their parametric knowledge. For example, LLMs tend to determine causal relationships based on the topological ordering of events (i.e., earlier events cause later ones), resulting in lower performance whenever events are not narrated in their exact causal order. Similarly, we demonstrate that LLMs struggle with long-term causal reasoning and often fail when the narratives are long and contain many events. Additionally, we show LLMs appear to rely heavily on their parametric knowledge at the expense of reasoning over the provided narrative. This degrades their abilities whenever the narrative opposes parametric knowledge. We extensively validate these failure modes through carefully controlled synthetic experiments, as well as evaluations on real-world narratives. Finally, we observe that explicitly generating a causal graph generally improves performance while naive chain-of-thought is ineffective. Collectively, our results distill precise failure modes of current state-of-the-art models and can pave the way for future techniques to enhance causal reasoning in LLMs.
@@ 1 Introduction
Causal reasoning is a core component of intelligence and decision-making, enabling an agent to move beyond associations (i.e., events that are observed to occur together) to determine the consequences of their actions (i.e., interventions) and answering what-if questions (counterfactuals). Causal models have been widely studied in machine learning and artificial intelligence (Peters et al., 2017; Pearl, 2009) as well as in the context of human cognition (Waldmann, 2017). The advent of Large Language Models (LLMs) has led to opportunities for leveraging large-scale textual data for improving causal inference (Liu et al., 2024; Zhang et al., 2023).
While many works have shown that LLMs are capable of memorizing and recalling causal knowledge, they can fail at reliably leveraging that knowledge for reasoning (Jin et al., 2023; Zečević et al., 2023). When reading a news article or a story, understanding why a particular event occurred requires reasoning over the causal structure of the events implied by the given narrative. Similarly, determining which events in the narrative will be affected by certain actions requires reasoning about which events are causally downstream of that action. In such cases, simply recalling parametric causal knowledge (that some event typically causes another event) may not suffice as the agent must leverage the causal structure inherent to the specific narrative at hand.
@@
In this work, we aim to understand the causal reasoning abilities of LLMs from textual narratives. We consider settings where there is an (unknown) underlying causal chain graph of the form $V_{1}\rightarrow V_{2}\rightarrow\ldots\rightarrow V_{N}$, where each node $V_{i}$ has some semantic meaning (e.g., smoking), that is verbalized in the form of a (realistic) narrative $N$. For a given narrative $N$, we consider two causal reasoning tasks: (1) Does $V_{i}$ have a causal effect on $V_{j}$? and (2) Given the node identities $(V_{1},\ldots V_{N})$, construct a causal chain graph faithful to the narrative. While these tasks do not encompass all aspects of causal reasoning (e.g., counterfactual reasoning), they are important primitives for successful causal reasoning.
@@
Our primary contribution is to shed light on the limitations and unreliable shortcuts that LLMs use for these two causal reasoning tasks. We focus on the following failure modes. Firstly, we show that LLMs heavily rely on the order in which the causal relationships are verbalized in the narrative. We observe that when the narrative is constructed in the reverse topological order of the causal chain (i.e., the edge $V_{i}\rightarrow V_{i+1}$ is narrated before $V_{i-1}\rightarrow V_{i}$), the performance of the LLM suffers as it often assigns the cause to an earlier event and the effect to a later event in the narrative. Secondly, we show that LLMs use their parametric causal knowledge (i.e., if an event typically causes another event) as a shortcut to answer causal questions. Thus, when the cause-and-effect pairs implied by the narrative conflict with the parametric knowledge, the LLM often ignores the specifics of the narrative and defaults to its parametric knowledge. This occurs even when the narratives are written to explicitly disagree with parametric knowledge and the LLM is prompted to ignore outside information. Thirdly, we find that LLMs fail more often as the narratives become longer.
Perhaps surprisingly, we also find that the LLM’s answers to cause-and-effect queries, including with Chain of Thought (CoT) (Wei et al., 2022) and In-Context Learning, do not have a high degree of consistency with the causal graph that the LLM extracts from the narrative. On the other hand, we observe that just using the extracted graph to make decisions often seems to alleviate the above failure modes. This finding is notable as it suggests that the specific formatting of the causal graph extraction task can more reliably elicit the language model’s causal reasoning capabilities relative to the baseline of CoT prompting.
We first validate our findings through a series of carefully controlled experiments  [^2] The code and data is at https://github.com/shantanu95/llm_causal_reasoning with synthetic causal events and narratives (Sec. 3). We use the LLM to generate a causal chain of events that form the ground truth causal graph. We then prompt the LLM to generate a narrative-based on that causal graph. Next, we conduct experiments on narratives generated from real-world causal chain graphs (Sec. 4) extracted from CauseNet (Heindorf et al., 2020), a dataset of (claimed) causal relationships between real-world concepts. We generate narratives for these chain graphs using the LLM as well as from sentences in CauseNet, ensuring that the narrative as a whole remains coherent.
Leveraging the LLM to verbalize causal relationships leads to more natural-sounding narratives (which is difficult with formulaic templates). Depending on the specific causal relationship, causality might be indicated using a phrase like ‘causes’, ‘leads to’, ‘resulted in’, etc. We focus on simple chain graphs to isolate the simplest cases of the failure modes, demonstrating that the LLMs take unreliable shortcuts on these simple graphs, both when the narratives are generated by the LLM itself as well as narratives with real-world sentences.
Taken together, our work makes the following contributions:
  1. 1.
    We demonstrate that even state-of-the-art large language models are unreliable at extracting causal relationships from realistic narratives expressing simple chain graphs.
  2. 2.
    Focusing on two key aspects of causal reasoning, we distill concrete failure modes to account for this unreliability: reliance on positional shortcuts, parametric knowledge, and narrative length.
  3. 3.
    We investigate the impact of various prompting strategies on reasoning ability, finding that incorporating the estimated causal graph structure is most helpful.
@@ 2 Related Works
Reasoning in Large Language Models
Recently, there has been significant interest in evaluating and improving the reasoning capabilities of large language models. Prior works have examined reasoning in diverse settings including mathematics (Cobbe et al., 2021; Hendrycks et al., 2021), social interactions (Sap et al., 2019), and common-sense tasks (Zellers et al., 2019). Through this extensive investigation, large language models have demonstrated unreliable reasoning capabilities, often failing unexpectedly on relatively straightforward queries (Wan et al., 2024; Nezhurina et al., 2024). Extensive prior work has also studied methods for improving the reasoning capabilities of large language models such as by eliciting step-by-step explanations (Wei et al., 2023), finetuning on reasoning traces (Zelikman et al., 2022), and training models to find and correct errors (Kumar et al., 2024). In this work, we perform a comprehensive study of the causal reasoning capabilities of large-language models, identifying the existence of key failure modes in state-of-the-art models.
@@
Prior works have also examined the causal reasoning capabilities of large-language models. Jin et al. (2023) develop a benchmark for testing causal reasoning in LLMs given causal graphs, finding that language models can struggle with the task. However, the queries examined in Jin et al. (2023) require probability calculations, potentially conflating causal reasoning and arithmetic failures. Zečević et al. (2023) hypothesizes that large language model may only be “imitating” causal reasoning abilities from their pretraining corpus. Their work additionally investigates what they call “natural word chains,” or how well the LLM is able to accurately identify causation of events when they are linked together in chain graphs. Our paper builds on this by performing various tests on narratives that correspond to graphs. Kıcıman et al. (2024) shows that LLMs have strong abilities to generate causal texts. We make use of this generation for narrative creation in our synthetic experiments. Tan et al. (2022) shows the capability of a neural network trained on news data to label causal structures in individual sentences.
Parametric Knowledge Conflicts
Large language models have been shown to memorize factual knowledge present in their pretraining corpora, resulting in parametric knowledge (Petroni et al., 2019; Jiang et al., 2020). Prior works have observed that language models can often fail to use information provided in their context as a result of such memorized knowledge (Krishna et al., 2021; Longpre et al., 2022). This has been identified as a significant hurdle in the reliability of summarization and retrieval augmented-based systems (Rehman et al., 2023; Jin et al., 2024). Xie et al. (2024) find that the coherence of the provided context can control whether a model chooses to use its parametric knowledge over contextual information. In this work, we examine the distinct setting of causal reasoning and demonstrate that it can also be harmed by over-reliance on parametric knowledge.
@@ 3 Experiments with Synthetic Data
3.1 Data Generation Process
In our synthetic experiments, we use two leading LLMs: OpenAI’s GPT-4o (OpenAI et al., 2024) and Anthropic’s Claude 3.5 Sonnet (Anthropic, 2024). The purpose of our synthetic setup is to carefully control the conditions under which the LLMs are tested. In terms of the general setup of our fully synthetic experiments, as we summarize in Figure 1, we first use the LLM to generate events (which are real world phenomena like rain or plants growing). Then these events are linked together into a chain graph $G$ that acts as the causal ground truth. The LLM is given $G$ and asked to create a narrative that stays faithful to the causal relationships in $G$.
Providing only the narrative as input (and not $G$), we then ask the LLM to find $G^{\prime}$, the prediction of the underlying causal structure expressed by the narrative. Next, a series of causal questions is created by randomly sampling $10$ tuples of events from $G$ and asking the LLM whether an event in the tuple causes the other based on the narrative and/or $G^{\prime}$. We also explore many mediums of prompting including Chain of Thought (COT) (Wei et al., 2023) and In-Context Learning (Dong et al., 2024) (see Appendix A for the prompts).
Figure 1: Summary of Synthetic Experiment Setup
@@ 3.1 Data Generation Process
3.2 Reliance on Topological Ordering
Our experiments show that LLMs rely on the ordering in which the events are verbalized in a narrative when determining causal relationships. To investigate this, we started with randomly generated events that were used to make a ground truth graph $G$. During the creation of the narrative, we specified that the LLM either place the events in (1) the order that matches the topological causal ordering of the graph (e.g., if event $A$ causes $B$, then event $A$ is mentioned before $B$ in the narrative), or (2) a way that runs opposite to the causal ordering (event $B$ would be mentioned before event $A$ in the narrative even though $A$ causes $B$). We refer to these as the Forward and a Reverse topological ordering, respectively. As an example, the following is a GPT-4o generated Reverse topological narrative for the causal chain: Art $\rightarrow$ Wine tasting $\rightarrow$ Charity fundraiser:
@@ 3.2 Reliance on Topological Ordering
Each edge in the narrative is verbalized in the opposite order to its place in the causal chain. All narratives can be found in the linked code. Our synthetic narratives are structured in a simple way such that any causal relations would be evident to a human, yet even despite that we will show stark failures of the LLM.
As shown in Figure 2 (left), in the Forward direction, a simple question prompting strategy asking causal questions (line labeled Forward Mean), naive CoT prompting (line labeled Forward-Chain Mean) where we ask the LLM to think step-by-step, and an In-Context Learning prompt (line labeled Forward-Context Mean) where we provide example narratives and causal answering all perform very well. This is in contrast to the Reverse orientation when we look at the performance of the simple prompt (line labeled Reverse Mean), naive CoT prompt (line labeled Reverse-Chain Mean), and In-Context Learning prompt (line labeled Reverse-Context Mean). From this plot, we can note that naive COT prompting and In-Context Learning prompts do not seem to significantly boost accuracy under our conditions. Perhaps more interestingly, we find that the way the LLM answers questions using the narratives is not always consistent with the causal graph $G^{\prime}$ that the LLM builds when asked to predict the underlying graph structure (see consistency plot in bottom of Figure 2, where consistency measures agreement between the answers of the LLM and $G^{\prime}$). We see that in the Reverse orientation, answers given by the extracted causal graph G’ and the previously discussed prompting strategies seem to differ greatly. Additionally, the trend of those prompting strategies on the consistency plot for the Forward orientation narratives (comparing performance to $G^{\prime}$) mirrored their trend on the accuracy plot which compares performance to ground truth $G$ (left side).
This led us to test the accuracy of only using the extracted graph $G^{\prime}$ to answer causal questions. In this case, once $G^{\prime}$ is extracted by the LLM it is not given to the LLM again to answer questions (but rather used directly). We found that this strategy did significantly better in the Reverse direction than the other prompting strategies. Surprisingly, using $G^{\prime}$ in the Reverse direction narratives to answer causal questions did as well as using $G^{\prime}$ in the Forward direction narratives.
Next, we tried prompting using the narrative and $G^{\prime}$ (the LLM is given $G^{\prime}$ in this case in the prompt). This technique could be thought of as a type of CoT prompting strategy. However, in the Reverse direction narratives, the increase in accuracy achieved by only using $G^{\prime}$ completely dissipates. We conjecture that the process of building the extracted Causal Graph $G^{\prime}$ forces the LLM to engage in long term reasoning instead of using the simple shortcut, but when the narrative is again provided - the LLM defaults its shortcut.
In the Claude 3.5 Sonnet (Fig. 4), The LLM also does far better with the Forward topological ordering than it does for the Reverse ordering. For the Reverse ordering, failure can again be mitigated almost completely by using only $G^{\prime}$ to answer questions.
@@
Figure 4: Claude 3.5 Sonnet Experiments: Left side plot mirrors the left side plot of Fig. 2 and the right side mirrors the plot in Fig. 3. Descriptions of prompt strategies match the before-mentioned figures.
3.3 Parametric Knowledge Conflict

### v3

@@ Abstract
The ability to robustly identify causal relationships is essential for autonomous decision-making and adaptation to novel scenarios. However, accurately inferring causal structure requires integrating both world knowledge and abstract logical reasoning. In this work, we investigate the interaction between these two capabilities through the representative task of causal reasoning over narratives. Through controlled synthetic, semi-synthetic and real-world experiments, we find that state-of-the-art large language models (LLMs) often rely on superficial heuristics—for example, inferring causality from event order or recalling memorized world knowledge without attending to context. Furthermore, we show that simple reformulations of the task can elicit more robust reasoning behavior. Our evaluation spans a range of causal structures, from linear chains to complex graphs involving colliders and forks. These findings uncover systematic patterns in how LLMs perform causal reasoning and lay the groundwork for developing methods that better align LLM behavior with principled causal inference.
@@ 1 Introduction
Large Language Models (LLMs) have achieved promising performance across a range of tasks, as a result of their ability to absorb rich world knowledge from large-scale unsupervised data. As such, there is rising interest in deploying them in autonomous and agentic settings. However, a key prerequisite for autonomous decision making is causal reasoning: agents must be able to move beyond relying on associations and be capable of robustly determining the consequences of their actions. Does this capability arise naturally from large-scale pretraining?
Robust causal reasoning is particularly challenging as it relies on a combination of knowledge and reasoning capabilities. Unlike mathematical reasoning benchmarks (Cobbe et al., 2021) – which draw from a relatively constrained set of problem solving strategies and results – arriving at correct causal inferences often requires leveraging domain-specific knowledge about the events or variables involved. On the other hand, models must go beyond blindly retrieving memorized associations and knowledge to identify the correct relationships under atypical or counter-intuitive settings. World knowledge and reasoning must be carefully balanced for robust causal reasoning: incorporating some memorized world knowledge can be helpful, but over-reliance can lead to errors in atypical settings.
Prior works have primarily studied reasoning and world-knowledge of LLMs separately. For example, benchmarks on mathematical reasoning or coding typically study these capabilities in isolation – with minimal external world knowledge needed to solve problems. On the other hand, benchmarks for knowledge intensive tasks can generally be solved by simple retrieval of memorized knowledge. Thus, the interplay of knowledge retrieval and reasoning (and potential conflicts between them) remains understudied.
@@
In this work, we study this interplay in the setting of causal reasoning over textual narratives. We pay particular attention to (a) long sequences of events (as are likely to arise in autonomous LLM settings) and (b) atypical settings where typical “common-sense" causal relationships may not hold. Concretely, we start by considering settings where there is an (unknown) underlying causal chain graph of the form $V_{1}\rightarrow V_{2}\rightarrow\ldots\rightarrow V_{N}$, where each node $V_{i}$ has some semantic meaning (e.g., smoking), that is verbalized in the form of a (realistic) narrative $N$. For a given narrative $N$, we consider two causal reasoning tasks: (1) Does $V_{i}$ have a causal effect on $V_{j}$? and (2) Given the node identities $(V_{1},\ldots V_{N})$, construct a causal chain graph that is faithful to the narrative. While these tasks do not encompass all aspects of causal reasoning (e.g., counterfactual reasoning), they are important primitives for successful causal reasoning. We then extend our analysis to more complex underlying graph structures including colliders and forks.
@@
Our primary contribution is to characterize when large language models are able to robustly reason under these causal structures. On the one hand, we find that LLMs display distinctive failure modes related to interference between reasoning and world knowledge in causal inference. Firstly, we show that LLMs are influenced heavily by a prior that causes are likely to appear before effects in a narrative. We observe that when the narrative is constructed in the reverse topological order of the causal chain (i.e., the edge $V_{i}\rightarrow V_{i+1}$ is narrated before $V_{i-1}\rightarrow V_{i}$), the performance of the LLM suffers as it often assigns the cause to an earlier event and the effect to a later event in the narrative. Secondly, we show that LLMs use their parametric causal knowledge (i.e., if an event typically causes another event) as a shortcut to answer causal questions. Thus, when the cause-and-effect pairs implied by the narrative conflict with the parametric knowledge, the LLM often ignores the specifics of the narrative and defaults to its parametric knowledge. Neither prompting with Chain of Thought (CoT) (Wei et al., 2022) nor In-Context Learning alleviates these failures.
However, LLMs are much less impacted by variation in the reasoning difficulty of the task when the prompting scheme explicitly isolates reasoning and world knowledge. First, we find that asking the LLM to extract the entire causal graph implied by the narrative results in a high degree of success at correctly ordering individual events, largely avoiding both failure modes described above. However, these benefits dissipate if the model is prompted to use the extracted graph alongside the narrative. Second, LLMs exhibit only slight performance degradation when reasoning over narratives that display more complex graph structures than chains, for example forks or colliders. Third, while LLMs often struggle with longer narratives containing more events, this failure is also substantially mitigated by asking the LLM to just extract a graph. All together, our results paint a more nuanced picture of LLMs’ causal capabilities than simple success or failure and suggest that future development should focus on isolating and then composing LLMs’ strengths at reasoning and world knowledge in order to avoid conflicts between them. We validate our findings across both controlled synthetic (LLM generated) narratives and real-world articles from CauseNet Heindorf et al. (2020).
Ultimately, our work presents a comprehensive study of causal reasoning in LLMs. Our findings reveal that current state-of-the-art models can fail to reliably perform simple causal reasoning tasks on narratives. We distill two shortcuts that underlie this unreliability: over reliance on positional biases and parametric knowledge. On the other hand, we find that incorporating an explicit causal graph identification task can elicit significantly better reasoning behavior.
@@ 2 Related Works
Jin et al. (2023) develop a benchmark for testing causal reasoning in LLMs given causal graphs, finding that language models can struggle with the task. However, the queries examined in Jin et al. (2023) require probability calculations, potentially conflating causal reasoning and arithmetic failures. Tan et al. (2022) shows the capability of a neural network trained on news data to label causal structures in individual sentences. Joshi et al. (2024a) chronicles failure modes in textual, but non-narrative form data (e.g. text formulaically written as Event 1 Causes Event 2 Causes Event 3 Causes Event 4). Our paper expands upon such a line of work by testing the LLM’s abilities in both real and synthetic texts that would plausibly be seen in everyday life. Another contrasting work, Jin et al. (2024), uses only statistical language indicating event correlations as input.
(Gordon et al., 2012; Joshi et al., 2024b; Ho et al., 2023; Zhang et al., 2023; Wang et al., 2023; Ashwani et al., 2024) study causal reasoning ability as it relates to inferring causal relations based on "common sense". In such common-sense based settings, it is straightforward for models to simply rely on memorized knowledge from pretraining and achieve good performance, without leveraging any more general causal reasoning capabilities. Our work seeks to disentangle this general causal reasoning ability by specifically testing cases where causal relationships may contradict common-sense knowledge. This serves as a more robust measurement of the causal reasoning capabilities in unfamiliar and atypical scenarios. Empirically, we show that models struggle significantly in adapting to unfamiliar causal relations.
Another important distinction of our work is the focus on longer-form narratives. Existing works such as (Gordon et al., 2012; Zečević et al., 2023; Ho et al., 2023; Frohberg and Binder, 2022; Li et al., 2023; Gao et al., 2023) primarily examine short-form questions about a single causal relationship. On the other hand, our work examines longer and more complex sequences of events. Moreover, in contrast to domain-specific question banks such as Intuitive Physics studied in (Zečević et al., 2023), our narratives examine a more diverse range of topics (as illustrated by the sample narratives presented). As a result, our dataset provides a more realistic and diverse examination of LLM causal reasoning capabilities than prior works. As such, our work is unique in that we are the first paper to analyze non-common sense based causal reasoning in narratives that use everyday language.
Causal Story Generation
Kıcıman et al. (2024) shows that LLMs have strong abilities to generate causal texts. Ammanabrolu et al. (2020) introduces soft causal relations—causal constraints that match what readers expect—and uses commonsense inferences to bridge high-level plot points, resulting in more coherent narratives that align with everyday causal expectations. Tian et al. (2021) contributes by employing counterfactual knowledge to generate hyperboles, making story generation more realistic. Li et al. (2022) shows that asking a model to explain a cause or effect by generating new text conflates language generation with prediction; instead, their approach asks the model to simply indicate the sentence number representing the cause or effect, leading to stories that better respect causal relations. For our synthetic text generation, we focus on creating narratives that are extremely explicit and simple. In contrast to Ammanabrolu et al. (2020) and Li et al. (2022)’s approach of bridging events using commonsense, our narrative scheme already embeds explicit causal language between events that are causally related so that no inference or common-sense reasoning is required from the reader to reason about causality. Furthermore, our experiments often intentionally contradict common-sense parametric knowledge to check the model’s ability to solely rely on the self-contained narrative. Similarly, regarding Tian et al. (2021), we opted to avoid abstract language structures that might confuse even human readers.
@@ 3 Experiments with Synthetic Data
3.1 Setting
Synthetic Narrative Generation
In our synthetic experiments, we use three leading LLMs: OpenAI’s GPT-4o (OpenAI et al., 2024), Anthropic’s Claude 3.5 Sonnet (Anthropic, 2024), and the open source LLama 3.1 8b (Grattafiori et al., 2024). While we focus on GPT-4o in the main text, results from other models are in the Appendix. The purpose of our synthetic setup is to carefully control the conditions under which the LLMs are tested. In terms of the general setup of our fully synthetic experiments, we first use the LLM to generate events (which are real world phenomena like rain or plants growing). Then these events are linked together into a chain graph $G$ that acts as the causal ground truth (eg rain $\rightarrow$ plants growing). The LLM is given $G$ and asked to create a narrative that stays faithful to the causal relationships in $G$. These narratives are checked by researchers to ensure consistency with their base causal graphs. More specifically, when constructing the dataset, we asked researchers (select authors who were blind to the true underlying graph) to reconstruct the causal chains given just the narratives, and 98 percent of the time (out of 100 random samples), the humans were able to find the unique correct causal ordering. Roughly 2500 narrative samples were generated. To ensure a variety of events go into the narratives, we generate 100 to 1000 distinct events at a time and randomly pick the small number needed for narrative construction (all narratives are in supplementary files and select narratives are in the Appendix).
Providing only the narrative as input (and not $G$), we then ask the LLM to find $G^{\prime}$, the predicted underlying causal structure expressed by the narrative. In other words, the LLM is asked to output a causal graph that it thinks embodies the relationships in the narrative. Next, a series of causal questions is created by randomly sampling $10$ tuples of events from $G$ and asking the LLM whether an event in the tuple causes the other based on the narrative and/or $G^{\prime}$.
Prompting Strategies
We evaluate five prompting styles for causal reasoning where the names in italics represent those used in the legends of figures: Standard QA Prompting (Standard), where the model is simply asked to identify the causal relation between two narrative events; Chain-of-Thought (CoT), which instructs the model to articulate step-by-step reasoning before answering; In-Context Learning (In-Context), which precedes the query with illustrative question–answer examples; Explicit Causal Graph Extraction (Graph), which asks the model to generate an entire causal graph $G^{\prime}$ over all events and assesses whether the ordering of the target pair is correct; Narrative-Augmented Graph Extraction (Narr-Graph), which first elicits $G^{\prime}$ and then supplies both $G^{\prime}$ and the original narrative for joint reasoning about the causal pair. Exact prompts are in Appendix A.
@@ 3.1 Setting
3.2 Impact of Event Ordering
Our experiments show that LLMs rely on the ordering in which the events are verbalized in a narrative when determining causal relationships. To investigate this, we started with randomly generated events that were used to make a ground truth graph $G$. During the creation of the narrative, we specified that the LLM either places the events in (1) the order that matches the topological causal ordering of the graph (e.g., if event $A$ (indirectly or directly) causes $B$, then event $A$ is mentioned before $B$ in the narrative), or (2) a way that runs opposite to the causal ordering (event $B$ would be mentioned before event $A$ in the narrative even though $A$ (directly or indirectly) causes $B$). We refer to these as the Forward and Reverse topological ordering, respectively. As an example, the following is a GPT-4o generated Reverse topological narrative for the causal chain: Art $\rightarrow$ Wine tasting $\rightarrow$ Charity fundraiser:
@@ 3.2 Impact of Event Ordering
Each edge in the narrative is verbalized in the opposite order to its place in the causal chain. All narratives can be found in the linked code.
LLMs Rely on Event Ordering Across Prompting Strategies
As shown in Figure 1 (left), in the Forward direction, standard QA, CoT, and In-Context prompts all perform very well. This is in contrast to the Reverse orientation when we look at the performance of the standard QA, COT, and In-Context prompts. From this plot, we can see that naive COT and In-Context prompting do not seem to significantly boost accuracy under our conditions. Perhaps more interestingly, we find that the way the LLM answers questions using the narratives is not always consistent with the causal graph $G^{\prime}$ that the LLM builds when asked to predict the underlying graph structure (see consistency plot in right side of Figure 1, where consistency measures agreement between the answers of the LLM and $G^{\prime}$). In the Reverse orientation, answers given by the extracted causal graph G’ and the previously discussed prompting strategies seem to differ greatly. Additionally, the trend of those prompting strategies on the consistency plot for the Forward orientation narratives (comparing performance to $G^{\prime}$) mirrors their trend on the accuracy plot which compares performance to ground truth $G$ (left side).
Explicit Causal Graph Extraction Avoids Shortcuts
This led us to test the accuracy of only using the extracted graph $G^{\prime}$ to answer causal questions. In this case, once $G^{\prime}$ is extracted by the LLM, it is not given to the LLM again to answer questions (but rather used directly). We found that this strategy did significantly better in the Reverse direction than the other prompting strategies. Surprisingly, using $G^{\prime}$ in the Reverse direction narratives to answer causal questions did as well as using $G^{\prime}$ in the Forward direction narratives. Next, we tried prompting using the narrative and $G^{\prime}$ (the LLM is given $G^{\prime}$ in this case in the prompt). This technique could be thought of as a type of CoT prompting strategy. However, in the Reverse direction narratives, the increase in accuracy achieved by only using $G^{\prime}$ completely dissipates. We conjecture that the process of building the extracted Causal Graph $G^{\prime}$ forces the LLM to engage in long term reasoning instead of using the simple shortcut, but when the narrative is again provided - the LLM defaults back to the shortcut.
@@
3.3 Impact of Parametric Knowledge (In)consistency
Experimental Setup